# Reducing Circuit Depth in Lindblad Simulation via Step-Size Extrapolation

### v2

@@ Appendix F Local error expansion of the dilated Hamiltonian approximation
    $\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\left[\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{l^{p+1}}{(p+1)!}\sum_{j=0}^{k-p}c_{0,j,k-p}\,s^{j}+\sum_{p=1}^{k}\frac{1}{(p+1)!}\sum_{j=0}^{k-p}c_{p+1,j,k-p}\,s^{j}\right]ds$ (152)
@@
    $\sum_{j=0}^{k}c_{0,j,k}\ t^{j}=\sum_{j=0}^{k}\left(\delta_{j,1}\cdot\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{l^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!}\right)t^{j}$ (153)
@@
    $\|\Gamma_{k}^{(i)}(t)\|\leq\sum_{j=0}^{k}l\cdot c_{i-1,j,k}\ t^{j}+\delta_{j,0}\cdot\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{l^{p+1}}{(p+1)!}\sum_{j=0}^{k-p}c_{i-1,j,k-p}t^{j}+\frac{\sum_{j=0}^{k-p}c_{i+p,j,k-p}t^{j}}{(p+1)!}$ (154)
@@
    $\sum_{j=0}^{k}c_{i,j,k}\ t^{j}=$ $lc_{i-1,k,k}\,t^{k}$ (155)
    $+$ $\sum_{j=0}^{k-1}\left(\,l\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{l^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!}\right)t^{j}\,$ (156)

### v3

@@ Appendix F Local error expansion of the dilated Hamiltonian approximation
    $\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\left[\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{\ell^{p+1}}{(p+1)!}\sum_{j=0}^{k-p}c_{0,j,k-p}\,s^{j}+\sum_{p=1}^{k}\frac{1}{(p+1)!}\sum_{j=0}^{k-p}c_{p+1,j,k-p}\,s^{j}\right]ds$ (152)
@@
    $\sum_{j=0}^{k}c_{0,j,k}\ t^{j}=\sum_{j=0}^{k}\left(\delta_{j,1}\cdot\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{\ell^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!}\right)t^{j}$ (153)
@@
    $\|\Gamma_{k}^{(i)}(t)\|\leq\sum_{j=0}^{k}\ell\cdot c_{i-1,j,k}\ t^{j}+\delta_{j,0}\cdot\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{\ell^{p+1}}{(p+1)!}\sum_{j=0}^{k-p}c_{i-1,j,k-p}t^{j}+\frac{\sum_{j=0}^{k-p}c_{i+p,j,k-p}t^{j}}{(p+1)!}$ (154)
@@
    $\sum_{j=0}^{k}c_{i,j,k}\ t^{j}=$ $\ell c_{i-1,k,k}\,t^{k}$ (155)
    $+$ $\sum_{j=0}^{k-1}\left(\,\ell\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{\ell^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!}\right)t^{j}\,$ (156)